### v1

# Relative topological principality and the ideal intersection property for groupoid C*-algebras

Abstract.
We introduce the notion of relative topological principality for a family $\{H_{\alpha}\}$ of open subgroupoids of a Hausdorff étale groupoid $G$. The C*-algebras $C^{*}_{r}(H_{\alpha})$ of the groupoids $H_{\alpha}$ embed in $C^{*}_{r}(G)$ and we show that if $G$ is topologically principal relative to $\{H_{\alpha}\}$ then a representation of $C^{*}_{r}(G)$ is faithful if and only if its restriction to each of the subalgebras $C^{*}_{r}(H_{\alpha})$ is faithful. This variant of the ideal intersection property potentially involves several subalgebras, and gives a new method of verifying injectivity of representations of reduced groupoid C*-algebras. As applications we prove a uniqueness theorem for Toeplitz C*-algebras of left cancellative small categories that generalizes a recent result of Laca and Sehnem for Toeplitz algebras of group-embeddable monoids, and we also discuss and compare concrete examples arising from integer arithmetic.
${}^{1}$ Supported by NSERC Discovery Grant RGPIN-2021-02459
${}^{2}$ Supported by NSERC Discovery Grant RGPIN-2017-04052
1. Introduction
One often wants a method for determining the injectivity of ***-homorphisms with domain a fixed C*-algebra $A$. In some situations there may be a C*-subalgebra $B$ of $A$ such that any ***-homomorphism is injective on $A$ if and only if it is injective on $B$; this is particularly useful if the subalgebra $B$ is more easily understood than the whole algebra $A$. Equivalently, we say that a C*-subalgebra $B$ of $A$ detects ideals in $A$, or has the ideal intersection property in $A$, if the only (closed 2-sided) ideal $J$ of $A$ such that $B\cap J=\{0\}$ is the zero ideal. If $B$ detects ideals in $A$, then we can verify injectivity of a ***-homomorphism on $A$ by checking that it is injective on $B$. There are many cases known where certain types of C*-algebras have important subalgebras with the ideal intersection property, e.g. [KawamuraTomiyama, ArchboldSpielberg, ExelLacaQuigg, exelNonHausdorffEtaleGroupoids2011, LS, brown2015cartan, Starling, kennedyIdealIntersectionProperty2021] (see Section 3 below).
In this paper we consider the ideal intersection property for subalgebras and, more generally, families of subalgebras of $C^{*}_{r}(G)$, where $G$ is a second-countable Hausdorff étale groupoid. We introduce the notion of a groupoid being relatively topologically principal with respect to a family of open subgroupoids. Our main result is that if $G$ is relatively topologically principal with respect to a family $\mathscr{H}$ of open subgroupoids then a representation of $C^{*}_{r}(G)$ is faithful if and only if it is faithful on $C^{*}_{r}(H)$ for each $H\in\mathscr{H}$. Our notion of relative topological principality is a generalization of effectiveness or topological principality for étale groupoids, but our main result also captures several cases of the ideal intersection property that had been previously studied outside of the context of effective groupoids, including results from [brown2015cartan], [LS], and [Starling].
The paper is structured as follows. Section 2 consists of preliminaries on inverse semigroups, étale groupoids, and the C*-algebras associated to these objects. In Section 3 we review several previous results about the ideal intersection property in order to set the context for the following section. In Section 4 we give the definition of relative topological principality and prove our main result, Theorem 4.6. We conclude by illustrating our result in two kinds of examples: In Section 5 we consider the C*-algebras associated by Spielberg to left cancellative small categories [SpielbergSmallCats], and in Section 6 we discuss concrete examples from integer arithmetic introduced in [Laca-Schulz23].
Acknowledgements: The results of this paper form part of the the second named author’s MSc Thesis [GG-MSCthesis2023] written under the supervision of the first and third named authors. We would like to thank Jack Spielberg for several helpful comments.
2. Preliminaries
By an étale groupoid $G$ we mean a topological groupoid with locally compact Hausdorff unit space, such that the source and target maps $\textbf{s},\textbf{t}:G\rightrightarrows G^{(0)}$ are local homeomorphisms. A bisection is a subset $A\subseteq G$ such that the source and target maps restrict to homeomorphisms $A\xrightarrow{\cong}\textbf{s}(A)$ and $A\xrightarrow{\cong}\textbf{t}(A)$ respectively. Recall that an inverse semigroup is a semigroup $S$ such that for each $s\in S$ there is a unique element $s^{-1}$, the inverse of $s$, satisfying the identities $ss^{-1}s=s$ and $s^{-1}ss^{-1}=s^{-1}$. The collection of open bisections of an étale groupoid $G$ is denoted $\operatorname{Bis}(G)$ and forms an inverse semigroup under elementwise multiplication and inversion of subsets:
$U\cdot V:=\{\alpha\beta\;|\;\alpha\in U,\beta\in V,(\alpha,\beta)\in G^{(2)}\},\;\;\;\;U^{-1}=\{\alpha^{-1}\;|\;\alpha\in U\}.$
Given a unit $x\in G^{(0)}$ the isotropy at $x$ is the group $G^{x}_{x}=\{\alpha\in G\;|\;\textbf{s}(\alpha)=x=\textbf{t}(\alpha)\}$. The isotropy of $G$ is the subgroupoid $\operatorname{Iso}(G):=\bigcup_{x\in G^{(0)}}G^{x}_{x}$. We denote by $\operatorname{Iso}(G)^{\circ}$ its topological interior.
We use the standard groupoid notation for more general small categories as well. Accordingly, if $\mathcal{C}$ is a small category, we write $\mathcal{C}_{x}^{y}$ to denote the set of morphisms from $x$ to $y$. The set of objects of $\mathcal{C}$ is denoted $\mathcal{C}^{(0)}$ and is identified with the collection of identities in the set of arrows $\mathcal{C}^{(1)}$. We typically suppress the superscript on the set of arrows and just write $\mathcal{C}=\mathcal{C}^{(1)}$. If $a\in\mathcal{C}_{y}^{z}$ and $b\in\mathcal{C}_{x}^{y}$ we write the composition of $a$ and $b$ as $ab$.
3. Ideal property and uniqueness theorems
An important early result in the study of the ideal intersection property appears in the work of Kawamura and Tomiyama [KawamuraTomiyama], where they show that when a discrete group $G$ acts on a space $X$, if the diagonal subalgebra $C(X)$ in the crossed product $C(X)\rtimes G$ has the ideal intersection property in $C(X)\rtimes G$ then the interior of the set of fixed points of $X$ for each nontrivial element of $G$ is empty; moreover, if $G$ is amenable then the converse holds. In [ArchboldSpielberg] Archbold and Spielberg generalize this result to the setting of discrete C*-dynamical systems where $C(X)$ is replaced by a not necessarily commutative C*-algebra $A$. They introduced the notion of a topologically free action, which is the appropriate notion of freeness in this setting, and showed that topological freeness of the action implies that any ideal $I\subseteq A\rtimes G$ that intersects trivially with $A$ must be contained in the kernel of the canonical quotient map $\lambda:A\rtimes G\rightarrow A\rtimes_{r}G$ from the full crossed product to the reduced one. In particular, when $G$ is amenable (so that $\lambda$ is injective), $A$ has the ideal intersection property in $A\rtimes G$. The result of Archbold and Spielberg indicates that if one wants to verify injectivity concretely, then one should look at the reduced crossed product so as not to be concerned with the ideal $\ker\lambda$.
Assisted by this intuition, Exel, Laca, and Quigg [ExelLacaQuigg] generalize the result of Kawamura and Tomiyama on a different axis than that of Archbold and Spielberg. They show that if a discrete group $G$ has a topologically free partial action on a space $X$ then $C_{0}(X)$ has the ideal intersection property in the reduced crossed product $C_{0}(X)\rtimes_{r}G$. Under a countability assumption this result generalizes nicely to C*-algebras of Hausdorff étale groupoids [exelNonHausdorffEtaleGroupoids2011]. For general étale groupoids the concept of topological freeness splits into several non-equivalent notions; these agree if the groupoid is second-countable and Hausdorff. Perhaps the most commonly used condition in the second-countable Hausdorff setting is the property of being topologically principal or effective which says that the interior of the isotropy is just the unit space.
One thing all the above examples have in common is that they primarily study the ideal intersection property for an inclusion $B\subseteq A$ where $B$ is the unit fiber of a Fell bundle that satisfies some sort of topological freeness condition. The first step away from the topologically free setting occurs in [brown2015cartan, Theorem 3.1], where Brown, Nagy, Reznikoff, Sims, and Williams show that if $G$ is a second-countable Hausdorff étale groupoid, which they do not require to be topologically principal (i.e. one may have $G^{(0)}\subsetneq\operatorname{Iso}(G)^{\circ}$), then $C^{*}_{r}(\operatorname{Iso}(G)^{\circ})$ has the ideal intersection property in $C^{*}_{r}(G)$. This reduces to the previous case in the event $G$ is topologically principal.
The theorem of Brown, Nagy, Reznikoff, Sims, and Williams represents a conceptual step forward for studying ideals in $C^{*}_{r}(G)$ in the abstract, but using it in concrete situations is not necessarily easy. Moreover, there are situations where one has an open subgroupoid $H$ of an étale groupoid $G$ that does not contain $\operatorname{Iso}(G)^{\circ}$ yet $C^{*}_{r}(H)$ still detects ideals in $C^{*}_{r}(G)$.
A very simple example of this is provided by the transformation groupoid associated to the action of a countable discrete group $G$ on itself by translation. The action is free so the interior of the isotropy is simply the unit space, hence the faithfulness criterion from [ArchboldSpielberg, brown2015cartan] says that a representation of $C_{0}(G)\rtimes_{r}G$ is faithful if and only if it is faithful on $C_{0}(G)$. But this does not take into account that the action is also minimal, and hence $C_{0}(G)\rtimes_{r}G\cong\mathcal{K}(\ell^{2}(G))$ is simple. With this in mind, we may simply take a single point, say, $e\in G$, for which the isotropy is $\{e\}\subset G$, and use instead the C*-algebra of the singleton subgroupoid $\{e\}\ltimes\{e\}$, namely $\mathbb{C}\delta_{e}$, which is much smaller than $C_{0}(G)$ yet is equally effective at detecting ideals in $C_{0}(G)\rtimes_{r}G$ (admittedly because there is only one nontrivial ideal to detect in this setting).
Less obvious examples have been studied using several different approaches. For instance, in [quasilat] it is first shown that the diagonal subalgebra has the ideal intersection property, but then faithfulness is further reduced to the nonvanishing of a collection of projections [quasilat, Theorem 3.7]. If the quasi-lattice semigroup is finitely generated, cf. [nica, Section 6.3], then a single projection, hence a $1$-dimensional subalgebra, does the job. A second, more recent example is the uniqueness theorem [LS, Theorem 5.1] valid for the reduced Toeplitz C*-algebra $\mathcal{T}_{r}(P)$ of a group embeddable monoid $P\subset G$. One can write $\mathcal{T}_{r}(P)$ as the C*-algebra of a partial transformation groupoid $G\ltimes\Omega$. The group $P^{*}$ of invertible elements of $P$ is a subgroup of $G$ and the transformation subgroupoid $P^{*}\ltimes\Omega$ embeds as a clopen subgroupoid of $G\ltimes\Omega$. According to [LS, Theorem 5.1], $C^{*}_{r}(P^{*}\ltimes\Omega)$ detects ideals in $C^{*}_{r}(G\ltimes\Omega)$. In many cases $P^{*}\ltimes\Omega$ is strictly smaller than $(\operatorname{Iso}G\ltimes\Omega)^{\circ}$.
In hindsight, the results described above can be reformulated as situations in which $C^{*}_{r}(H)$ detects ideals in $C^{*}_{r}(G)$, where $H$ is an open subgroupoid of $G$. In particular, $C^{*}_{r}(H)$ has the ideal intersection property in $C^{*}_{r}(G)$ in the following cases:
  • $G$ is the action groupoid of a discrete group acting (or partially acting) topologically freely on a locally compact Hausdorff space and $H$ is the unit space [ExelLacaQuigg] [KawamuraTomiyama].
  • $G$ is a topologically principal second-countable Hausdorff étale groupoid and $H$ is the unit space [exelNonHausdorffEtaleGroupoids2011].
  • $G$ is any second-countable Hausdorff étale groupoid and $H=\operatorname{Iso}(G)^{\circ}$ [brown2015cartan].
  • G is the partial transformation groupoid associated to a group embeddable monoid $P$ and $H$ is the transformation groupoid obtained by restricting the action to the group of invertible elements of $P$ [LS].
Motivated by this we aim to find a condition on open subgroupoids of a groupoid $G$ that guarantees that when taken together, their reduced C*-algebras detect ideals in the reduced C*-algebra of $G$.
4. Relative topological principality
With an eye on exploiting the orbit structure in the unit space in order to simplify ideal detection, we now introduce our main concept.
Definition 4.1.
Let $G$ be an étale groupoid and let $\mathscr{H}$ be a family of open subgroupoids. We say that $G$ is topologically principal relative to $\mathscr{H}$ if the set
$X_{\mathscr{H}}:=\{u\in G^{(0)}:\exists\Gamma\in\operatorname{Bis}(G)\text{ and }H\in\mathscr{H}\text{ s.t. }u\in\Gamma^{-1}\Gamma\text{ and }\Gamma G^{u}_{u}\Gamma^{-1}\subseteq H\}$ (4.2)
is dense in $G^{(0)}$.
It will be convenient to have a reciprocal terminology; when $G$ is topologically free relative to a family $\mathscr{H}$ of open subgroupoids we say that $\mathscr{H}$ constrains the isotropy of $G$. If the family is a singleton =$\mathscr{H}=\{H\}$, as will often be the case, we will simply say that $G$ is topologically principal relative to $H$, or that $H$ constrains the isotropy of $G$, and write $X_{H}:=X_{\{H\}}$.
Before we state our main result, Theorem 4.6 below, we give a few examples that illustrate the notion of relative topological principality, especially when framed within the various contexts discussed in Section 3.
Example 4.3.
Let $G$ be a countable discrete group acting on itself and consider the trivial subgroupoid $H=\{(e,e)\}$ of the transformation groupoid $G\ltimes G$. Then $X_{H}=G$, hence $G\ltimes G$ is topologically principal relative to $H$.
Example 4.4.
Let $G$ be an étale groupoid and let $H:=G^{(0)}$ be the unit space, viewed as an open subgroupoid. If $\Gamma G^{u}_{u}\Gamma^{-1}\subseteq G^{(0)}$ for some open bisection $\Gamma$ then $\Gamma G^{u}_{u}\Gamma^{-1}$ is the trivial group, and so is $G^{u}_{u}$. By (4.2) $u\in X_{H}$ if and only if $G^{u}_{u}=\{u\}$, that is, $X_{H}$ is the set of units with trivial isotropy. As a consequence, $G$ is topologically principal if and only if it is topologically principal relative to the unit space.
Example 4.5.
Let $G$ be a second-countable Hausdorff étale groupoid. By [brown2015cartan, Lemma 3.3] the set $\{u\in G^{(0)}\;|\;G^{u}_{u}\subseteq\operatorname{Iso}(G)^{\circ}\}$ is dense in $G^{(0)}$ hence $G$ is always topologically principal relative to $\operatorname{Iso}(G)^{\circ}$.
Theorem 4.6.
Let $G$ be a second-countable Hausdorff étale groupoid and suppose that $G$ is topologically principal relative to a family $\mathscr{H}$ of open subgroupoids . Then a representation $\rho$ of $C^{*}_{r}(G)$ is faithful if and only if its restriction to $C^{*}_{r}(H)\subseteq C^{*}_{r}(G)$ is faithful for each $H\in\mathscr{H}$.
To prove Theorem 4.6 we will need to use isomorphisms of C*-algebras of subgroupoids that are inner in an appropriate sense. It is natural that in a groupoid each element $\gamma\in G$ will act as a partial symmetry, but strictly speaking, one can only conjugate $\alpha\in G$ by $\gamma$ if $\textbf{s}(\gamma)=\textbf{t}(\alpha)$ and $\textbf{s}(\alpha)=\textbf{t}(\gamma^{-1})$, in which case $\textbf{s}(\alpha)=\textbf{t}(\alpha)$, so $\alpha$ must be in the isotropy bundle of $G$. This will not be sufficient for our purposes so we broaden our understanding of what it means for $G$ to act on itself by conjugation by allowing an action of $\operatorname{Bis}(G)$ on the category Sub$(G)$ of subgroupoids of $G$. If $G$ is a group then $\operatorname{Bis}(G)$ may essentially be identified with $G$ and one may recover the inner automorphisms of $G$ from this action.
We formalize this notion in the following definition.
Definition 4.7.
By a partial automorphism of an étale groupoid $G$ we mean an isomorphism $H\rightarrow K$ between two subgroupoids of $G$ in the category of topological groupoids. We say that a partial automorphism $\phi$ is inner if there is an open bisection $\Gamma\in\operatorname{Bis}(G)$ such that $\phi(h)=\Gamma h\Gamma^{-1}$ for all $h\in H$. Here $\Gamma h\Gamma^{-1}:=(\textbf{s}|_{\Gamma}^{-1}\circ\textbf{t})(h)\cdot h\cdot(\textbf{t}|_{\Gamma^{-1}}^{-1}\circ\textbf{s})(h)$.
Lemma 4.8.
Let $G$ be an étale groupoid, $\Gamma\in\operatorname{Bis}(G)$, and $H$ a subgroupoid of $G$ such that $H^{(0)}\subseteq\textbf{s}(\Gamma)$. Then the map $\operatorname{Ad}_{\Gamma}:H\rightarrow\Gamma H\Gamma^{-1}\text{ given by }\alpha\mapsto\Gamma\alpha\Gamma^{-1}:=(\textbf{s}|_{\Gamma}^{-1}\circ\textbf{t})(\alpha)\cdot\alpha\cdot(\textbf{t}|_{\Gamma^{-1}}^{-1}\circ\textbf{s})(\alpha)$ is an inner partial automorphism of $G$.
Proof.
Let $(\alpha,\beta)\in H^{(2)}$. Since $\Gamma$ is a bisection and $H^{(0)}\subseteq\textbf{s}(\Gamma)$ there are unique elements $\gamma_{1},\gamma_{2},\gamma_{3}\in\Gamma$ such that
$\textbf{t}(\alpha)=\textbf{s}(\gamma_{1}),\quad\textbf{s}(\alpha)=\textbf{t}(\gamma_{2}^{-1})=\textbf{s}(\gamma_{2})=\textbf{t}(\beta),\quad\text{and }\textbf{s}(\beta)=\textbf{t}(\gamma_{3}^{-1}).$
Thus $\Gamma\alpha\Gamma^{-1}\Gamma\beta\Gamma^{-1}=\gamma_{1}\alpha\gamma_{2}^{-1}\gamma_{2}\beta\gamma_{3}^{-1}=\gamma_{1}\alpha\beta\gamma_{3}^{-1}=\Gamma\alpha\beta\Gamma^{-1}$ so $\operatorname{Ad}_{\Gamma}$ is a homomorphism. Continuity follows from continuity of the source, range, and multiplication maps. Since $\operatorname{Ad}_{\Gamma^{-1}}$ is a continuous inverse, $\operatorname{Ad}_{\Gamma}$ is an isomorphism of topological groupoids. ∎
Definition 4.9.
Let $B$ and $C$ be C*-subalgebras of a C*-algebra $A$. We say a subalgebra isomorphism $\phi:B\rightarrow C$ of $A$ is inner if there is a partial isometry $v$ in the universal enveloping von Neumann algebra $A^{**}$ of $A$ such that $\phi(b)=vbv^{*}$ and $\phi^{-1}(c)=v^{*}cv$ for all $b\in B$ and $c\in C$.
Remark 4.10.
We use the term “subalgebra isomorphism” to distinguish this notion from the more restrictive one of partial automorphism used in connection to partial actions of groups; specifically, partial automorphisms are subalgebra isomorphisms for which domain and range are 2-sided ideals. This feature is required only for the definition of a crossed product, and does not concern us here.
It is a standard theme that inner automorphisms preserve more structure than outer ones. For our approach to detecting ideals, the most important consequence of innerness is the following.
Lemma 4.11.
Let $B$ and $C$ be C*-subalgebras of a C*-algebra $A$. Suppose $\rho:A\rightarrow{\mathbf{B}}(\mathcal{H})$ is a representation and $\phi:B\rightarrow C$ is an inner subalgebra isomorphism of $A$. Then $\rho$ is faithful on $B$ if and only if $\rho$ is faithful on $C$.
Proof.
Suppose $\rho$ is faithful on $B$ and let $\tilde{\rho}$ denote the unique weak-* continuous extension to $A^{**}$. Let $c\in C$ be arbitrary and let $b\in B$ satisfy $c=vbv^{*}=\phi(b)$, where $v\in A^{**}$ is the partial isometry implementing $\phi$. Note that $\left\lVert b\right\rVert=\left\lVert c\right\rVert$ because $\phi$ is an isomorphism. Since $\rho$ is a ***-homomorphism $\left\lVert\rho(c)\right\rVert\leq\left\lVert c\right\rVert$. As for the reverse inequality we have
$\left\lVert c\right\rVert=\left\lVert b\right\rVert=\left\lVert\rho(b)\right\rVert=\left\lVert\rho(v^{*}vbv^{*}v)\right\rVert=\left\lVert\tilde{\rho}(v^{*})\rho(vbv^{*})\tilde{\rho}(v)\right\rVert\leq\left\lVert\rho(vbv^{*})\right\rVert=\left\lVert\rho(c)\right\rVert;$
thus $\rho$ is faithful on $C$. The converse follows by symmetry because $\phi^{-1}:C\rightarrow B$ is an inner subalgebra isomorphism. ∎
Next we adapt Lemma 3.7 and Proposition 3.8 from [kennedyIdealIntersectionProperty2021] and show that the the reduced C*-algebra of a groupoid $G$ has an action of $\operatorname{Bis(G)}$ by subalgebra isomorphisms of $C^{*}_{r}(G)$.
Lemma 4.12.
Suppose $G$ is a Hausdorff étale groupoid, and let $S$ denote the collection of partial isometries in the enveloping von Neumann algebra $C^{*}_{r}(G)^{**}$. Then there is a multiplicative and involutive map $v:\operatorname{Bis}(G)\rightarrow S$ mapping $\Gamma\mapsto v_{\Gamma}$ such that for every $\Gamma\in\operatorname{Bis}(G)$ and every open subgroupoid $H$ such that $H^{(0)}\subseteq\textbf{s}(\Gamma)$,
  1. (1)
    the net $(f)_{0\leq f\leq{\mathbf{1}}_{\Gamma}}\subseteq C_{c}(\Gamma)$ converges to $v_{\Gamma}$ in the weak-* topology as $f$ increases to ${\mathbf{1}}_{\Gamma}$.
  2. (2)
    $v_{\Gamma}^{*}v_{\Gamma}={\mathbf{1}}_{\Gamma^{-1}\Gamma}$ and $v_{\Gamma}v_{\Gamma}^{*}={\mathbf{1}}_{\Gamma\Gamma^{-1}}$.
  3. (3)
    If $g\in C_{c}(V)$ for some $V\in\operatorname{Bis}(H)$ then $v_{\Gamma}g=g(\Gamma^{-1}\cdot)\in C_{c}(\Gamma V)$, and $gv_{\Gamma}^{*}=g(\cdot\Gamma)\in C_{c}(V\Gamma^{-1})$.
  4. (4)
    The isomorphism $\phi:C^{*}_{r}(H)\rightarrow C^{*}_{r}(\Gamma H\Gamma^{-1})$ induced by the groupoid isomorphism $H\rightarrow\Gamma H\Gamma^{-1}$ is implemented by the conjugation $\operatorname{Ad}_{v_{\Gamma}}:a\mapsto v_{\Gamma}av_{\Gamma}^{*}$ and is therefore an inner subalgebra isomorphism.
Proof.
[…]
which is to say $(f)_{0\leq f\leq{\mathbf{1}}_{\Gamma}}\rightarrow{\mathbf{1}}_{\Gamma}$ in the weak-* topology on $B^{\infty}(\Gamma)$ and hence in $C^{*}_{r}(G)^{**}$ because $C_{0}(\Gamma)^{**}$ is weak-* closed in the latter. Setting $v_{\Gamma}:={\mathbf{1}}_{\Gamma}$ gives part (1). Next let $g\in C_{c}(V)$ for an open bisection $V$ of $H$. Then for all $0\leq f\leq{\mathbf{1}}_{\Gamma}$ we have
$(f*g)(x)=\begin{cases}f(y)g(y^{-1}x)&\text{if there exists a (necessarily unique) }y\in\Gamma\text{ such that }y^{-1}x\in V,\\ 0&\text{otherwise.}\end{cases}$
The convolution operator $\psi_{g}:B^{\infty}(\Gamma)\rightarrow B^{\infty}(\Gamma V)$ defined by $\psi_{g}(f)=f*g$ is continuous with respect to the topology of pointwise convergence on $B^{\infty}(\Gamma)$ and $B^{\infty}(\Gamma V)$, hence
$(v_{\Gamma}g)(x)=\lim_{0\leq f\leq{\mathbf{1}}_{\Gamma}}(f*g)(x)=g(\Gamma^{-1}x)$
and similarly
$(gv_{\Gamma^{-1}})(x)=\lim\limits_{0\leq f\leq{\mathbf{1}}_{\Gamma^{-1}}}(g*f)(x)=g(x\Gamma),$
proving part (3). To see that the map $\operatorname{Bis}(\Gamma)\rightarrow S$, $\Gamma\mapsto v_{\Gamma}$ is a semigroup homomorphism observe that
$v_{\Gamma_{1}}v_{\Gamma_{2}}=\lim_{0\leq f\leq{\mathbf{1}}_{\Gamma_{2}}}v_{\Gamma_{1}}f=\lim_{0\leq f\leq{\mathbf{1}}_{\Gamma_{2}}}f(\Gamma_{1}^{-1}\cdot)=\lim_{0\leq f\leq{\mathbf{1}}_{\Gamma_{1}\Gamma_{2}}}f=v_{\Gamma_{1}\Gamma_{2}}$
and similarly $v_{\Gamma}^{*}=v_{\Gamma^{-1}}$. Moreover, this gives part (2) because
$v_{\Gamma}v_{\Gamma}^{*}=\lim_{0\leq f\leq{\mathbf{1}}_{\Gamma^{-1}}}v_{\Gamma}f=\lim_{0\leq f\leq{\mathbf{1}}_{\Gamma^{-1}}}f(\Gamma^{-1}\cdot)={\mathbf{1}}_{\Gamma\Gamma^{-1}}$
and similarly $v_{\Gamma}^{*}v_{\Gamma}={\mathbf{1}}_{\Gamma^{-1}\Gamma}$.
Finally we prove part (4). By (3) we have $v_{\Gamma}gv_{\Gamma}^{*}=g(\Gamma^{-1}\cdot\Gamma)\in C_{c}(\Gamma H\Gamma^{-1})$ for each $g\in C_{c}(H)$. By hypothesis we have $\Gamma^{-1}\Gamma H\Gamma^{-1}\Gamma=H$ and thus $\operatorname{Ad_{v_{\Gamma}^{*}}}:C_{c}(\Gamma H\Gamma^{-1})\rightarrow C_{c}(H)$ is the inverse of $\operatorname{Ad}_{v_{\Gamma}}$. Since $\left\lVert a\right\rVert=\left\lVert v_{\Gamma}^{*}v_{\Gamma}av_{\Gamma}^{*}v_{\Gamma}\right\rVert\leq\left\lVert v_{\Gamma}av_{\Gamma}^{*}\right\rVert\leq\left\lVert a\right\rVert$, the result follows. ∎
Remark 4.13.
We point out that Lemma 4.12 is about the reduced groupoid C*-algebra $C^{*}_{r}(G)$ instead of the universal one considered in [kennedyIdealIntersectionProperty2021, Lemma 3.7.] and that we do not need to assume the unit space $G^{(0)}$ to be compact because we do not rely upon the notion of $G$-C*-algebra used in [kennedyIdealIntersectionProperty2021]. One should also notice that the set of partial isometries in a von Neumann algebra does not typically form an inverse semigroup under multiplication as claimed in [kennedyIdealIntersectionProperty2021, Lemma 3.7.]. This does not affect the results because all that is required is that the image of the map be an inverse semigroup.
The following two lemmas generalize results from [brown2015cartan] to a collection of subalgebras and subgroupoids.
Lemma 4.14 (cf. [brown2015cartan] Theorem 3.2. ).
Let $\mathscr{B}$ be a collection of C*-subalgebras of a C*-algebra $A$. For each $B\in\mathscr{B}$ suppose that $S_{B}$ is a collection of states on $B$ such that each state $\varphi\in S_{B}$ has a unique extension to a state $\tilde{\varphi}$ on $A$. Let $S:=\bigcup_{B\in\mathscr{B}}S_{B}$ and suppose that the sum of GNS representations $\bigoplus_{\varphi\in S}\pi_{\tilde{\varphi}}$ of the extensions of the states in $S$ is faithful on $A$. Then a representation $\rho:A\rightarrow{\mathbf{B}}(\mathcal{H})$ is faithful if and only if it is faithful on $B$ for each $B\in\mathscr{B}$.
Proof.
Suppose that the restriction of $\rho:A\rightarrow{\mathbf{B}}(\mathcal{H})$ to $B$ is faithful for each $B\in\mathscr{B}$. Set $J:=\operatorname{Ker}\rho$, let $B\in\mathscr{B}$ be arbitrary, and observe that $J\cap B=\{0\}$. By e.g. [dixmierAlgebras1982, Corollary 1.8.4.], $A_{0}:=J+B$ is a C*-subalgebra of $A$ with ideal $J$ and quotient $A_{0}/J\cong B$. Denote the quotient map by $\gamma:A_{0}\rightarrow B$. Let $\varphi\in S_{B}$ be arbitrary. Then $\tilde{\varphi}(a)=\varphi(\gamma(a))$ for all $a\in A_{0}$, for otherwise any extension of $\varphi\circ\gamma$ to $A$ would extend $\varphi$ and be different from $\tilde{\varphi}$.
Let $x\in J$ and denote by $\xi_{\tilde{\varphi}}$ the GNS vector of $\tilde{\varphi}$. For every $a\in A$,
$\langle\pi_{\tilde{\varphi}}(x^{*}x)\pi_{\tilde{\varphi}}(a)\xi_{\tilde{\varphi}},\pi_{\tilde{\varphi}}(a)\xi_{\tilde{\varphi}}\rangle=\langle\pi_{\tilde{\varphi}}(a^{*}x^{*}xa)\xi_{\tilde{\varphi}},\xi_{\tilde{\varphi}}\rangle=\tilde{\varphi}(a^{*}x^{*}xa)=\varphi(\gamma(a^{*}x^{*}xa))=0$
because $a^{*}x^{*}xa\in J$. Therefore $\pi_{\tilde{\varphi}}(x^{*}x)=0$, and since $\bigoplus_{\varphi\in S}\pi_{\tilde{\varphi}}$ is faithful, and $B\in\mathscr{B}$ and $\varphi\in S_{B}$ are arbitrary we must have $x=0$, proving that $\rho$ is faithful. ∎
Lemma 4.15 (cf. [brown2015cartan] Lemma 3.3.).
Let $G$ be a second-countable Hausdorff étale groupoid that is topologically principal relative to a family $\mathscr{H}$ of open subgroupoids. Let $X_{\mathscr{H}}$ be the set defined in (4.2) and suppose that $u\in X_{\mathscr{H}}$ and $\Gamma\in\operatorname{Bis}(G)$ satisfy $u\in\Gamma^{-1}\Gamma$ and $\Gamma G^{u}_{u}\Gamma^{-1}\subseteq H$ for some $H\in\mathscr{H}$. Then for each $f\in C_{c}(G)$ there exists $b\in C_{c}((\Gamma^{-1}H\Gamma)^{(0)})^{+}$ such that $\lVert b\rVert=b(u)=1$ and $bfb\in C_{c}(\Gamma^{-1}H\Gamma)$.
Proof.
Let $f\in C_{c}(G)$ and write $f=\sum_{D\in F}f_{D}$ where $F$ is a finite collection of precompact open bisections and $\operatorname{supp}(f_{D})\subseteq D$. For each $D\in F$ choose an open neighborhood $V_{D}$ of $u$ in $G^{(0)}$ as follows.
  • If there exists $\alpha\in D\cap G^{u}_{u}$ then $D\cap\Gamma^{-1}H\Gamma$ is an open neighborhood of $\alpha$, hence $V_{D}:=\textbf{s}(D\cap\Gamma^{-1}H\Gamma)\cap\textbf{t}(D\cap\Gamma^{-1}H\Gamma)$ is an open neighborhood of $u$ and $V_{D}DV_{D}\subseteq D\cap\Gamma^{-1}H\Gamma$.
  • If $D\cap G^{u}_{u}=\emptyset$ but there exists $\alpha\in D$ with $r(\alpha)=u\neq s(\alpha)$ or $r(\alpha)\neq u=s(\alpha)$; then take an open neighborhood $U\subseteq D$ of $\alpha$ such that $s(U)\cap r(U)=\emptyset$ and set $V_{D}:=r(U)$ so that $V_{D}DV_{D}=\emptyset$.
  • If $u\not\in r(D)\cup s(D)$ then take an open neighborhood $V_{D}$ of $u$ such that $f_{D}|_{V_{D}DV_{D}}=0$.
Let $V:=\cap_{D\in F}V_{D}$. Then $V$ is open and contains $u$. Choose $b\in C_{c}(V)^{+}$ such that $b(u)=\left\lVert b\right\rVert=1$. By construction $\operatorname{supp}(bfb)\subseteq\Gamma^{-1}H\Gamma$ as desired. ∎
Lemma 4.16 (cf. [brown2015cartan] Lemma 3.5).
Under the same assumptions as in Lemma 4.15, let $a\in C^{*}_{r}(G)$ and $\epsilon>0$. Then there exists $c\in C^{*}_{r}(\Gamma^{-1}H\Gamma)$ and $b\in C_{c}(G^{(0)})^{+}$ such that $\left\lVert b\right\rVert=1$ and $\varphi(b)=1$ for all states $\varphi$ that factor through $C_{r}^{*}(G_{u}^{u})$ and such that $\left\lVert bab-c\right\rVert<\epsilon$.
Proof.
Let $\epsilon>0$ and $a\in C^{*}_{r}(G)$. Let $f\in C_{c}(G)$ such that $\left\lVert a-f\right\rVert<\epsilon$. Apply Lemma 4.15 to find $b\in C_{c}(G^{(0)})^{+}$ such that $\left\lVert b\right\rVert=b(u)=1$ and $bfb\in C_{c}(\Gamma^{-1}H\Gamma)$. Let $\vartheta_{u}:C^{*}_{r}(\Gamma^{-1}H\Gamma)\rightarrow C^{*}_{r}(G^{u}_{u})$ be the completely positive contraction that extends the restriction map $C_{c}(\Gamma^{-1}H\Gamma)\rightarrow C_{c}(G^{u}_{u}),\ f\mapsto f|_{G^{u}_{u}}$ by [christensenNonExoticCompletions2022, Lemma 1.2]. Then $\vartheta_{u}(b)=\delta_{u}$ is the identity of $C^{*}_{r}(G^{u}_{u})$ hence if $\varphi$ is a state on $C^{*}_{r}(\Gamma^{-1}H\Gamma)$ such that $\varphi=\psi\circ\vartheta_{u}$ for some state $\psi$ on $C^{*}_{r}(G^{u}_{u})$ we have $\varphi(b)=\psi(\delta_{u})=1$. Moreover if we set $c=bfb$ we have
$\left\lVert bab-c\right\rVert=\left\lVert bab-bfb\right\rVert<\epsilon$
as desired. ∎
Before we prove our main theorem we need one final lemma.
Lemma 4.17.
Suppose that $G$ is a Hausdorff étale groupoid and let $x\in G^{(0)}$. Let
  • $\Phi:C^{*}_{r}(G)\rightarrow C_{0}(G^{(0)})$ be the canonical faithful conditional expectation,
  • $\vartheta_{x}:C^{*}_{r}(G)\rightarrow C^{*}_{r}(G^{x}_{x})$ be the completely positive contraction extending the restriction map $C_{c}(G)\rightarrow C_{c}(G^{x}_{x})$ as in [christensenNonExoticCompletions2022, Lemma 1.2.],
  • $\tau:C^{*}_{r}(G^{x}_{x})\rightarrow\mathbb{C}$ be the canonical faithful trace given by $\tau(a)=\langle\delta_{x},a\delta_{x}\rangle$ and
  • $\operatorname{ev}_{x}:C_{0}(G^{(0)})\rightarrow\mathbb{C}$ be evaluation at $x$.
Then the following diagram commutes.
[…]
Proof.
If $f\in C_{c}(G)$ then $\tau(\vartheta_{x}(f))=\langle\delta_{x},f|_{G^{x}_{x}}\delta_{x}\rangle=f(x)=\operatorname{ev}_{x}\circ\Phi$. The result follows by continuity. ∎
We are now ready for the proof of our main result.
Proof of Theorem 4.6.
For each $u\in X_{\mathscr{H}}$, $H\in\mathscr{H}$, and $\Gamma\in\operatorname{Bis}(G)$ such that $u\in\Gamma^{-1}\Gamma$ and $\Gamma G_{u}^{u}\Gamma^{-1}\subseteq H$ let $S_{u,\Gamma,H}$ denote the collection of states on $C^{*}_{r}(\Gamma^{-1}H\Gamma)$ that factor through $C^{*}_{r}(G^{u}_{u})$. Let $\rho:C^{*}_{r}(G)\rightarrow{\mathbf{B}}(\mathcal{H})$ be a representation that is faithful on the family of C*-subalgebras $\{C^{*}_{r}(H)\;|\;H\in\mathscr{H}\}$. Then $\rho$ is faithful on $C^{*}_{r}(K)\subseteq C^{*}_{r}(H)$ for every open subgroupoid $K$ of $H$. Let $K:=\textbf{t}(\Gamma)H\textbf{t}(\Gamma)$. Then $K\subseteq H$ is an open subgroupoid, with $K^{(0)}\subseteq\textbf{s}(\Gamma)$, and moreover $\Gamma^{-1}K\Gamma=\Gamma^{-1}H\Gamma$. By Lemma 4.12 $\operatorname{Ad}_{v_{\Gamma}^{*}}:C^{*}_{r}(K)\rightarrow C^{*}_{r}(\Gamma^{-1}H\Gamma)$ is an inner subalgebra isomorphism. Hence by Lemma 4.11 faithfulness of $\rho$ on $C^{*}_{r}(H)$ implies faithfulness on $C^{*}_{r}(\Gamma^{-1}H\Gamma)$.
Thanks to Lemma 4.16 we may apply [ArmstrongUniqueness, Theorem 5.4.], and conclude that for each $u\in X_{\mathscr{H}}$, $H\in\mathscr{H}$ and $\Gamma$ with $H^{(0)}\subseteq\textbf{s}(\Gamma)$ such that $G^{u}_{u}\subseteq\Gamma^{-1}H\Gamma$, every state $\varphi\in S_{u,\Gamma,H}$ has a unique state extension $\tilde{\varphi}$ to $C^{*}_{r}(G)$. By Lemma 4.14, in order to prove the theorem it suffices to prove that the direct sum $\pi_{S_{u,\Gamma,H}}:=\bigoplus_{\varphi\in S_{u,\Gamma,H}}\pi_{\tilde{\varphi}}$ of the GNS representations of the extended states is faithful on $C^{*}_{r}(G)$.
In order to prove this, suppose $\pi_{S_{u,\Gamma,H}}(a)=0$ and consider the canonical faithful conditional expectation $\Phi_{G}:C^{*}_{r}(G)\rightarrow C_{0}(G^{(0)})$.
We claim that $\Phi_{G}(a^{*}a)=0$, which then forces $a=0$ because $\Phi_{G}$ is faitfhul. Suppose for contradiction that $\Phi_{G}(a^{*}a)\neq 0$. Then $\Phi_{G}(a^{*}a)$ is a nontrivial positive function on $G^{(0)}$ so by density of $X_{\mathscr{H}}$ in $G^{(0)}$ there is a $u\in X_{\mathscr{H}}$ such that $\Phi_{G}(a^{*}a)(u)>0$. Let $\epsilon=\frac{\Phi_{G}(a^{*}a)(u)}{2}$. By Lemma 4.16 we can find $b\in C_{c}(G^{(0)})^{+}$ with $\|b\|=1$ and $c\in C^{*}_{r}(\Gamma^{-1}H\Gamma)$ such that $\left\lVert ba^{*}ab-c\right\rVert<\epsilon/2$ and $\varphi(b)=1$ for all states that factor through $C^{*}_{r}(G^{u}_{u})$. By assumption, $\pi_{S_{u,\Gamma,H}}(a)=0$ and hence also $\pi_{S_{u,\Gamma,H}}(ba^{*}ab)=0$ because $\left\lVert\pi_{S_{u,\Gamma,H}}(ba^{*}ab)\right\rVert\leq\left\lVert\pi_{S_{u,\Gamma,H}}(a^{*}a)\right\rVert=0$. If $\varphi\in S_{u,\Gamma,H}$ and $\{e_{\lambda}\}$ is an approximate unit in $C^{*}_{r}(G)$, then
$\tilde{\varphi}(ba^{*}ab)=\lim_{\lambda}\langle\pi_{\tilde{\varphi}}(ba^{*}ab)(e_{\lambda}+N_{\tilde{\varphi}}),(e_{\lambda}+N_{\tilde{\varphi}})\rangle=0$
and therefore
$|\tilde{\varphi}(c)|=|\tilde{\varphi}(ba^{*}ab-c)|\leq\left\lVert ba^{*}ab-c\right\rVert<\epsilon/2.$
Let $\vartheta_{u}:C^{*}_{r}(G)\rightarrow C^{*}_{r}(G^{u}_{u})$ be the completely positive contraction extending the restriction map. Then
$\left\lVert\vartheta_{u}(c)\right\rVert=\sup\big{\{}|\psi(\vartheta_{u}(c))|\;|\;\psi\text{ is a state on }C^{*}_{r}(G^{u}_{u})\big{\}}\leq\sup\big{\{}|\varphi(c)|\;|\;\varphi\in S_{u,\Gamma,H}\big{\}}\leq\epsilon/2.$
Now let $\Phi_{\Gamma,H}:C^{*}_{r}(\Gamma^{-1}H\Gamma)\rightarrow C_{0}((\Gamma^{-1}H\Gamma)^{0})$ be the conditional expectation given by restricting functions to the unit space. For $x\in(\Gamma^{-1}H\Gamma)^{0}$ let $\operatorname{ev}_{x}$ be evaluation at $x$ on $C_{0}((\Gamma^{-1}H\Gamma)^{0})$. Then $\operatorname{ev}_{x}\circ\Phi_{\Gamma,H}$ is a state on $C^{*}_{r}(\Gamma^{-1}H\Gamma)$; moreover for each $x\in X_{\Gamma^{-1}H\Gamma}=\{x\in(\Gamma^{-1}H\Gamma)^{0}\;|\;G_{x}^{x}\subseteq\Gamma^{-1}H\Gamma\}$ the state $\operatorname{ev}_{x}\circ\Phi_{\Gamma,H}$ factors through $C^{*}_{r}(G^{x}_{x})$ by Lemma 4.17. Let $\tau:C^{*}_{r}(G^{u}_{u})\rightarrow\mathbb{C}$ be the canonical tracial state. Since $\Phi_{G}|_{\Gamma^{-1}H\Gamma}=\Phi_{\Gamma,H}$ we have
$|\operatorname{ev}_{u}\circ\Phi_{G}(c)|=|\operatorname{ev}_{u}\circ\Phi_{\Gamma,H}(c)|=|\tau(\vartheta_{u}(c))|\leq\left\lVert\vartheta_{u}(c)\right\rVert\leq\epsilon/2.$
Since $\Phi_{G}$ is $C_{0}(G^{(0)})$-bilinear and $\operatorname{ev}_{u}$ is multiplicative,
$\operatorname{ev}_{u}\circ\Phi_{G}(ba^{*}ab)=\operatorname{ev}_{u}(b)(\operatorname{ev}_{u}\circ\Phi_{G})(a^{*}a)\operatorname{ev}_{u}(b)=(\operatorname{ev}_{u}\circ\Phi_{G})(a^{*}a).$
Thus
$|\operatorname{ev}_{u}\circ\Phi_{G}(a^{*}a)|$ $\leq|\operatorname{ev}_{u}\circ\Phi_{G}(a^{*}a)-\operatorname{ev}_{u}\circ\Phi_{G}(c)+\operatorname{ev}_{u}\circ\Phi_{G}(c)|$
$\leq|\operatorname{ev}_{u}\circ\Phi_{G}(ba^{*}ab)-\operatorname{ev}_{u}\circ\Phi_{G}(c)|+|\operatorname{ev}_{u}\circ\Phi_{G}(c)|$
$\leq\left\lVert ba^{*}ab-c\right\rVert+|\operatorname{ev}_{u}\circ\Phi_{G}(c)|$
$<\epsilon.$
But this contradicts our choice of $\epsilon$, hence $\Phi_{G}(a^{*}a)=0$ and therefore $a=0$. It follows that $\pi_{S_{u,\Gamma,H}}$ is injective. ∎
5. C*-algebras of Small Categories
In [SpielbergSmallCats] Spielberg associates certain C*-algebras to left cancellative small categories (LCSCs), generalizing his earlier work on C*-algebras of categories of paths [SpielbergPathCats]. These include C*-algebras of k-graphs and quasi-lattice ordered groups.
Definition 5.1.
A left cancellative small category (LCSC) is a small category $\mathcal{C}$ such that $xy=xz\implies y=z$ for all $x,y,z\in\mathcal{C}$ with $xy,xz\in\mathcal{C}$.
A left cancellative monoid is just a LCSC with one object, and the aim of this section is to use the results of the previous section to provide a version of the faithfulness/uniqueness theorem of Laca and Sehnem [LS, Theorem 5.1] in the more general setting of LCSCs.
The key fact is that left cancellation implies that the partially defined left multiplication maps given by $\textbf{s}(c)\mathcal{C}\rightarrow c\mathcal{C},\ x\mapsto cx$ for $c\in\mathcal{C}$ are partial bijections. Hence for each $c\in\mathcal{C}$ there is a partial isometry $T_{c}:\ell^{2}(\mathcal{C})\rightarrow\ell^{2}(\mathcal{C})$ determined by
$T_{c}\delta_{x}=\begin{cases}\delta_{cx}&\text{ if }\textbf{s}(c)=\textbf{t}(x)\\ 0&\text{otherwise}.\end{cases}$
The map $c\mapsto T_{c}$ is the left regular representation $T:\mathcal{C}\to{\mathbf{B}}(\ell^{2}(\mathcal{C}))$, and we are interested in the C*-algebra generated by its image.
Definition 5.2.
([LiGarsideCats, Definition 2.2]) The left reduced C*-algebra of the LCSC $\mathcal{C}$ is
$C^{*}_{r}(\mathcal{C}):=C^{*}(\{T_{c}\;|\;c\in\mathcal{C}\})\subseteq{\mathbf{B}}(\ell^{2}(\mathcal{C})).$
On our way towards Spielberg’s groupoid model for $C^{*}_{r}(\mathcal{C})$ we will use the inverse semigroup approach to $C^{*}_{r}(\mathcal{C})$ described in [LiGarsideCats] because it allows us to make easy comparison with semigroup C*-algebras. Thus we turn our attention to a natural inverse semigroup arising from a LCSC.
Definition 5.3.
([LiGarsideCats, Definition 2.3]) The left inverse hull $\mathcal{I}_{l}(\mathcal{C})$ of a LCSC $\mathcal{C}$ is the inverse subsemigroup of the symmetric monoid $\mathcal{I}(\mathcal{C})$ generated by the left multiplication maps $\textbf{s}(c)\mathcal{C}\rightarrow c\mathcal{C},\ x\mapsto cx$ for $c\in\mathcal{C}$.
When no confusion can arise we may omit $\mathcal{C}$ from the notation and simply write $\mathcal{I}_{l}$ instead of $\mathcal{I}_{l}(\mathcal{C})$. We often abuse notation and identify an element $c\in\mathcal{C}$ with the corresponding left multiplication map $c\in\mathcal{I}_{l}(\mathcal{C})$. We denote its inverse map in $\mathcal{I}_{l}$ by $c^{-1}$; by definition, this is the map defined by $c\mathcal{C}\rightarrow\textbf{s}(c)\mathcal{C},\ cx\mapsto x$, but when $c$ is invertible in $\mathcal{C}$, the inverse map $c^{-1}$ coincides with (left multiplication by) the element $c^{-1}\in\mathcal{C}$. Notice that the set $\mathcal{C}^{*}:=\mathcal{C}\cap\mathcal{C}^{-1}$ of invertibles in $\mathcal{C}$ is a groupoid.
An arbitrary element $s\in\mathcal{I}_{l}$ may be written in the form $d_{n}^{-1}c_{n}\cdots d_{1}^{-1}c_{1}$ where $d_{i},c_{i}\in\mathcal{C}$ for $i=1,2,\ldots,n$ and the successive concatenations are pairwise composable in $\mathcal{C}$; these are the zigzag maps of [SpielbergSmallCats, Definition 2.6]. The representation $T:\mathcal{C}\to{\mathbf{B}}(\ell^{2}(\mathcal{C}))$ can be extended to all of $\mathcal{I}_{l}$ by setting $T_{d_{n}^{-1}c_{n}\cdots d_{1}^{-1}c_{1}}:=T_{d_{n}}^{*}T_{c_{n}}\cdots T_{d_{1}}^{*}T_{c_{1}}$, see [SpielbergSmallCats, Lemma 11.1]. Clearly $C^{*}(\{T_{c}\;|\;c\in\mathcal{C}\})=\overline{\operatorname{span}}\{T_{s}\;|\;s\in\mathcal{I}_{l}\}$. So $C^{*}_{r}(\mathcal{C})$ is precisely the C*-algebra denoted by $\mathcal{T}_{\ell}(\mathcal{C})$ and called the regular Toeplitz algebra of $\mathcal{C}$ by Spielberg in [SpielbergSmallCats, Definition 11.2].
As pointed out at the beginning of [SpielbergSmallCats, Section 8], there are C*-algebraic relations satisfied in $C_{r}^{*}(\mathcal{C})=\mathcal{T}_{\ell}(\mathcal{C})$ that are not captured by the bare inverse semigroup structure of $\mathcal{I}_{l}(\mathcal{C})$, see also [LS, Section 3]. Indeed, in some cases the equality $f=\vee_{i=1}^{n}e_{i}$ holds for idempotents in $\mathcal{I}_{l}$ and hence is respected by the left regular representation, but it is not derived from the inverse semigroup structure alone. One needs to consider the order theoretic relations between idempotents. Recall that $E(\mathcal{I}(X))$ denotes the semilattice of idempotents of the symmetric monoid $\mathcal{I}(X)$ on a set $X$, and that $E(\mathcal{I}(X))$ is canonically isomorphic to the power set $\mathcal{P}(X)$, which is a lattice. We use this isomorphism to identify idempotents in $\mathcal{I}_{l}(\mathcal{C})$ with subsets of $\mathcal{C}$.
Remark 5.4.
The zero element in the symmetric monoid $\mathcal{I}(\mathcal{C})$ is the partial bijection whose domain and range are the empty set; it is also the zero of the semilattice of idempotents $E(\mathcal{I}(\mathcal{C}))\cong\mathcal{P(C)}$. We denote this element by 00. Whether $0\in\mathcal{I}_{l}$ depends on whether the empty set may be written as an intersection $\bigcap_{s\in F}\operatorname{dom}(s)$ for a finite set of elements $F\subseteq\mathcal{I}_{l}$.
The images of the idempotents in $\mathcal{I}_{l}$ form a commuting family of projections that is closed under multiplication, so their closed linear span is a commutative C*-subalgebra of $C^{*}_{r}(\mathcal{C})$.
Definition 5.5.
Let $\mathcal{C}$ be a LCSC. We define the diagonal subalgebra of $C^{*}_{r}(\mathcal{C})$ to be $D_{r}(\mathcal{C}):=\overline{\operatorname{span}}\{T_{e}\;|\;e\in E(\mathcal{I}_{l})\}.$
In order to lighten the notation, for the remainder of this section we make the following simplifications:
  • $\mathcal{C}$ is a LCSC
  • $\mathcal{I}_{l}=\mathcal{I}_{l}(\mathcal{C})$ is the left inverse hull of $\mathcal{C}$
  • $E=E(\mathcal{I}_{l})$ is the semilattice of idempotents of $\mathcal{I}_{l}$
  • $D_{r}=D_{r}(\mathcal{C})$ is the diagonal subalgebra of $C^{*}_{r}(\mathcal{C})$
  • $\Omega:=\{\chi\in\widehat{E}\;|\;\chi(f)=\vee_{i=1}^{n}\chi(e_{i}),\text{ whenever }f,e_{i}\in\mathcal{I}_{l}\text{ and }f=\vee_{i=1}^{n}e_{i}\}$ is the subspace of $\widehat{E}$ consisting of semicharacters preserving any finite suprema that exist in $\mathcal{I}_{l}$.
Next we show that the necessary extra relations on $E$ can be captured via the spectrum of $D_{r}$.
Lemma 5.6.
The map that sends a character $\chi\in\operatorname{Spec}(D_{r})$ to its restriction $\chi|_{E}$ to the subset $\{T_{e}\;|\;e\in E\}\subset D_{r}$ is a homeomorphism of $\operatorname{Spec}(D_{r})$ onto $\Omega$.
Proof.
The map $T:E\rightarrow D_{r},\ e\mapsto T_{e}$ is a faithful representation of the semigroup $E$. The restriction $\chi|_{E}$ of $\chi\in\operatorname{Spec}(D_{r})$ to this semigroup is a semicharacter, but it also preserves joins in $E$ because whenever $e,f\in E$ with $e\vee f\in E$ we have $T_{e\vee f}=T_{e}+T_{f}-T_{ef}$, which is respected by the multiplicative linear functional $\chi$. Conversely every $\chi\in\Omega$ preserves the multiplicative and linear relations among the $T_{e}$ hence extends to a character on $D_{r}=\overline{\operatorname{span}}\{T_{e}\;|\;e\in E\}$. This shows that restriction is a bijection between $\operatorname{Spec}(D_{r})$ and $\Omega$. To show that this bijection is indeed a homeomorphism we note that $D_{r}$ admits a surjection $C^{*}(E)\rightarrow D_{r}$ from the universal C*-algebra generated by a representation of the inverse semigroup $E$. Explicitly $C^{*}(E)$ is generated by projections $\{t_{e}\;|\;e\in E\}$ subject to the relations $t_{e}t_{f}=t_{ef}$ and $t_{e^{*}}=t_{e}^{*}$. This realizes $\operatorname{Spec}(D_{r})$ as a subspace of $\operatorname{Spec}(C^{*}(E))$. On the other hand the universal property of $C^{*}(E)$ gives a 1-1 correspondence between semicharacters $E\rightarrow\{0,1\}$ and characters $C^{*}(E)\rightarrow\mathbb{C}$. Since the weak-* topology on $\operatorname{Spec}(C^{*}(E))$ is the topology of pointwise convergence on $\widehat{E}$, and since the diagram
[…]
is commutative, the result follows because the bottom horizontal map can be seen as a restriction of the top horizontal map, which is a homeomorphism. ∎
Remark 5.7.
An alternate proof of Lemma 5.6 may be obtained using [cuntzKTheoryGroupAlgebras2017, Corollary 5.6.26.].
Just as in the case of semigroups, here too the spectrum of the diagonal can be seen as a compactification of $\mathcal{C}$ via a canonical dense embedding $\mathcal{C}\hookrightarrow\Omega$ as principal characters.
Lemma 5.8.
[LiGarsideCats, Lemma 2.12.] For each $c\in\mathcal{C}$ define $\chi_{c}:E\rightarrow\{0,1\}$ by
$\chi_{c}(e)=\begin{cases}1&\text{ if }cc^{-1}\leq e\\ 0&\text{ otherwise.}\end{cases}$
Then $\chi_{c}$ is a semicharacter and $\{\chi_{c}\;|\;c\in\mathcal{C}\}$ is a dense subset of $\Omega$.
Proof.
One easily verifies that each $\chi_{c}\in\Omega$. Recall that a basic open subset of the spectrum $\widehat{E}$ of the semilattice $E$ is of the form $\widehat{E}(e,F):=\{\chi\;|\;\chi(e)=1,\chi(f)=0\text{ for all }f\in F\}$. Density follows immediately from the observation that if $\widehat{E}(e,F)\neq\emptyset$ then $e\setminus\cup F\neq\emptyset$, in which case $\chi_{c}\in\widehat{E}(e,F)$ for each $c\in e\setminus\cup F$. ∎
Every inverse semigroup $S$ has a canonical action by partial homeomorphisms of the spectrum $\widehat{E(S)}$ of its semilattice of idempotents. The domain of $s\in S$ is $D_{s^{-1}s}=\{\chi\in\widehat{E(S)}\;|\;\chi(s^{-1}s)=1\}$ and the image under $s$ of $\chi\in D_{s^{-1}s}$ is the character $s\cdot\chi$, given by $(s\cdot\chi)(e)=\chi(s^{-1}es)$. It is easy to verify that $s\cdot\chi\in D_{ss^{-1}}$. We next show that $\Omega$ is invariant under this action.
Lemma 5.9.
$\Omega$ is an $\mathcal{I}_{l}$-invariant subset of $\widehat{E}$.
Proof.
Let $\chi\in\Omega$ and suppose $\chi\in D_{s^{-1}s}$ for some $s\in\mathcal{I}_{l}$. We claim that $s\cdot\chi\in\Omega$. Suppose $f$ is an idempotent which can be written as a join $f=\bigvee_{i=1}^{n}e_{i}$ of a finite set of idempotents. Then
$(s\cdot\chi)\left(\bigvee_{i=1}^{n}e_{i}\right)=\chi\left(s^{-1}\bigvee_{i=1}^{n}e_{i}s\right)=\chi\left(\bigvee_{i=1}^{n}s^{-1}e_{i}s\right)=\bigvee_{i=1}^{n}\chi(s^{-1}e_{i}s)=\bigvee_{i=1}^{n}(s\cdot\chi)(e_{i})$
so $\chi\in\Omega$ as desired. ∎
In order to obtain a groupoid that includes the extra relations encoded in the diagonal, we simply enlarge the underlying inverse semigroup in such a way that the idempotents become a Boolean algebra.
Definition 5.10.
(cf. [SpielbergSmallCats, Definition 2.8.]) Let $\mathcal{C}$ be a LCSC and let $E$ be the semilattice of idempotents of $\mathcal{I}_{l}$. The boolean closure $\overline{E}$ of $E$ in $\mathcal{I(C)}$ is the collection of subsets of $\mathcal{C}$ of the form $\bigvee_{i=1}^{n}\left(e_{i}\setminus\bigvee_{j=1}^{k_{i}}f_{ij}\right)$ where $f_{ij},e_{i}\in E$ and $f_{ij}\leq e_{i}$ for all $1\leq j\leq k_{i}$ and $1\leq i\leq n$. We also enlarge the left inverse hull $\mathcal{I}_{l}(\mathcal{C})$ to
¯$\overline{\mathcal{I}}_{l}(\mathcal{C}):=\{se\;|\;s\in\mathcal{I}_{l},e\in\overline{E},e\leq s^{-1}s\}$
which is an inverse semigroup with semilattice of idempotents $\overline{E}$.
It follows from Lemma 5.9 that the action of $\mathcal{I}_{l}$ on $\widehat{E}$ restricts to an action of $\mathcal{I}_{l}$ on $\Omega$. Now for every $\chi\in\Omega$ we may extend $\chi$ to an element $\overline{\chi}$ of $\widehat{\overline{E}}$ by setting $\overline{\chi}(\bigvee_{i=1}^{n}(e_{i}\setminus\bigvee_{j=1}^{k_{i}}f_{ij}))=\bigvee_{i=1}^{n}(\chi(e_{i})\setminus\bigvee_{j=1}^{k_{i}}\chi(f_{ij}))$, where the operations on the right side are to be understood in the sense of the two-element Boolean algebra $\{0,1\}$. This gives an embedding $\Omega\hookrightarrow\widehat{\overline{E}}$ and the canonical action of $\overline{\mathcal{I}_{l}}$ on $\widehat{\overline{E}}$ restricts to an action on $\Omega$.
Definition 5.11.
(cf. [LiGarsideCats, Remark 2.16]) Let ¯$\overline{\mathcal{I}}_{l}*\Omega=\{(s,\chi)\in\overline{\mathcal{I}}_{l}\times\Omega\;|\;\chi(s^{-1}s)=1\}$, and define an equivalence relation on ¯$\overline{\mathcal{I}}_{l}*\Omega$ by setting
$(s,\chi)\sim(t,\psi)\iff\begin{cases}\chi=\psi\text{ and}\\ \text{there is an idempotent }e\in\overline{\mathcal{I}}_{l}\text{ such that }\chi(e)=1\text{ and }se=te.\end{cases}$
We denote the equivalence class of a pair $(s,\chi)$ by $[s,\chi]$.
The transformation groupoid of a left cancellative small category $\mathcal{C}$ has underlying set ¯$\overline{\mathcal{I}}_{l}\ltimes\Omega:=(\overline{\mathcal{I}}_{l}*\Omega)/\sim$. The source and target maps are given by $\textbf{s}([r,\chi])=\chi$ and $\textbf{t}([r,\psi])=r\cdot\psi$ respectively. The inverse of $[s,\chi]\in\overline{\mathcal{I}}_{l}\ltimes\Omega$ is given by $[s^{-1},s\cdot\chi]$ and the product of $[s,t\cdot\chi]$ and $[t,\chi]$ is $[st,\chi]$. Then ¯$\overline{\mathcal{I}}_{l}\ltimes\Omega$ becomes an étale groupoid when given the topology with basis of open sets $\{[s,U]\;|\;s\in\overline{\mathcal{I}}_{l},U\text{ is an open subset of }D_{s^{-1}s}\}.$
In order to apply our relative topological principality results to the reduced C*-algebra of a LCSC we will rely on its realization as a groupoid C*-algebra given by Spielberg.
Theorem 5.12.
[SpielbergSmallCats, Theorem 11.6.] Let $\mathcal{C}$ be a LCSC and suppose that the groupoid ¯$\overline{\mathcal{I}}_{l}\ltimes\Omega$ is Hausdorff. Then $C^{*}_{r}(\mathcal{C})\cong C^{*}_{r}(\overline{\mathcal{I}}_{l}\ltimes\Omega)$.
For the remainder of the section we let $\mathscr{G}:=\overline{\mathcal{I}_{l}}\ltimes\Omega$. We need to understand the isotropy in $\mathscr{G}$ next. Recall that $\mathcal{C}^{*}=\mathcal{C}\cap\mathcal{C}^{-1}$ denotes the groupoid of invertible elements in $\mathcal{C}$ and note that $\mathscr{G}^{(0)}=(\overline{\mathcal{I}}_{l}\ltimes\Omega)^{(0)}=\Omega$. Our description of isotropy will occur in Lemma 5.16 below; we first need a few technical lemmas.
Lemma 5.13.
Suppose that $e\in\mathcal{C}^{(0)}$ and $u\in\mathcal{C}^{*}$ satisfy $u^{-1}u=e=uu^{-1}$. Then $[u,\chi_{e}]\in\mathscr{G}^{\chi_{e}}_{\chi_{e}}$.
Proof.
Suppose that $f\in E(\mathcal{I}_{l})$. Then
$\chi_{e}(f)=\begin{cases}1&\text{ if }e\leq f\\ 0&\text{otherwise,}\end{cases}$
so we need to show that $e\leq f\iff e\leq u^{-1}fu$. If $e\leq f$ i.e. $ef=e$ then $uu^{-1}f=uu^{-1}\implies u^{-1}uu^{-1}fu=u^{-1}uu^{-1}u=u^{-1}u=e$ that is $e(u^{-1}fu)=e$. Conversely if $e(u^{-1}fu)=e$ then $(u^{-1}u)(u^{-1}fu)=uu^{-1}$ so conjugating by $u$ we get $(uu^{-1})(uu^{-1})f(uu^{-1})=uu^{-1}$ which is equivalent to $ef=e$ since $e=uu^{-1}=u^{-1}u$ and idempotents commute. ∎
Lemma 5.14.
If $s\in\mathcal{I}_{l}$, $p,q\in\mathcal{C}$, and $p\in\operatorname{dom}(s)$, then $s(pq)=s(p)q$.
Proof.
If $s=d^{-1}c$ for $d,c\in\mathcal{C}$ then $cp=dr$ for some $r\in\mathcal{C}$ hence $cpq=drq$ and it follows that $s(pq)=rq=s(p)q$. We proceed by induction on the word length of $s=d_{k}^{-1}c_{k}\cdots d_{1}^{-1}c_{1}$. The desired result holds for $k=1$; assume it holds for all $k\leq\ell$. Let $p,pq\in\operatorname{dom}(s)$ where $s=d_{\ell+1}^{-1}c_{\ell+1}\cdots d_{1}^{-1}c_{1}$. Then $d_{1}^{-1}c_{1}(p)$ and $d_{1}^{-1}c_{1}(pq)=d_{1}^{-1}c_{1}(p)q$ are in $\operatorname{dom}(d_{\ell+1}^{-1}c_{\ell+1}\cdots d_{2}^{-1}c_{2})$. By induction we have $d_{\ell+1}^{-1}c_{\ell+1}\cdots d_{1}^{-1}c_{1}(pq)=d_{\ell+1}^{-1}c_{\ell+1}\cdots d_{2}^{-1}c_{2}(d_{1}^{-1}c_{1}(p)q)=d_{\ell+1}^{-1}c_{\ell+1}\cdots d_{1}^{-1}c_{1}(p)q$. ∎
Lemma 5.15.
Suppose that $e\in\mathcal{C}^{(0)}$ and $s\in\mathcal{I}_{l}$ satisfy $s^{-1}s=e=ss^{-1}$. Then $s\in\mathcal{C}^{*}$.
Proof.
Let $p\in e\mathcal{C}e$. Then, using Lemma 5.14, we have $s(p)=s(ep)=s(e)p$, so $s$ is left multiplication by $s(e)$. Similarly $s^{-1}$ is left multiplication by $s^{-1}(e)$; hence $s^{-1}(e)=s(e)^{-1}$ and $s\in\mathcal{C}^{*}$. ∎
We are now ready to describe the isotropy at the dense subset of principal characters in $\mathscr{G}^{(0)}$.
Lemma 5.16.
For each $c\in\mathcal{C}$, the isotropy at $\chi_{c}$ is $\mathscr{G}^{\chi_{c}}_{\chi_{c}}=\{[cuc^{-1},\chi_{c}]\;|\;u\in\textbf{s}(c)\mathcal{C}^{*}\textbf{s}(c)\}$.
Proof.
Let $e=\textbf{s}(c)$. Then
$\mathscr{G}_{\chi_{e}}^{\chi_{e}}=\{[u,\chi_{e}]\;|\;u\in\mathcal{I}_{l},u^{-1}u=e=uu^{-1}\}=\{[u,\chi_{e}]\;|\;u\in\mathcal{C}^{*},u^{-1}u=e=uu^{-1}\},$
where the first equality is from Lemma 5.13 and the second equality holds by Lemma 5.15. We then have $\mathscr{G}_{\chi_{c}}^{\chi_{c}}=[c,\chi_{e}]\mathscr{G}_{\chi_{e}}^{\chi_{e}}[c,\chi_{e}]^{-1}=\{[cuc^{-1},\chi_{c}]\;|\;u\in\textbf{s}(c)\mathcal{C}^{*}\textbf{s}(c)\}$. ∎
We use the above to show that the subgroupoid $\mathcal{C}^{*}\ltimes\Omega$ constrains the isotropy in $\mathscr{G}$.
Proposition 5.17.
The groupoid $\mathscr{G}=\overline{\mathcal{I}}_{l}\ltimes\Omega$ is topologically principal relative to $\mathcal{C}^{*}\ltimes\Omega$.
Proof.
Given $c\in\mathcal{C}$ the subset $[c,\Omega(c^{-1}c)]$ of $\mathscr{G}$ is an open bisection with $\textbf{s}([c,\Omega(c^{-1}c)])=\Omega(c^{-1}c)$ and $\textbf{t}([c,\Omega(c^{-1}c)])=\Omega(cc^{-1})$. Let $\Gamma:=[c,\Omega(c^{-1}c)]^{-1}$. Then $\Gamma\mathscr{G}^{\chi_{c}}_{\chi_{c}}\Gamma^{-1}=\mathscr{G}^{\chi_{\textbf{s}(c)}}_{\chi_{\textbf{s}(c)}}\subseteq\mathcal{C}^{*}\ltimes\Omega$; that is, the isotropy at each element of $X:=\{\chi_{c}\;|\;c\in\mathcal{C}\}$ may be conjugated into $\mathcal{C}^{*}\ltimes\Omega$ by an open bisection. Since $X$ is dense by Lemma 5.8, $\mathscr{G}$ is topologically principal relative to $\mathcal{C}^{*}\ltimes\Omega$. ∎
An application of Theorem 4.6 now yields a faithfulness criterion for representations of $C^{*}_{r}(\mathcal{C})$.
Theorem 5.18.
Let $\mathcal{C}$ be a left cancellative small category such that $\mathscr{G}:=\overline{\mathcal{I}_{l}}\ltimes\Omega$ is second-countable and Hausdorff. A representation of $C_{r}^{*}(\mathcal{C})\cong C^{*}_{r}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ is faithful if and only if its restriction to the C*-subalgebra $C^{*}_{r}(\mathcal{C}^{*}\ltimes\Omega)$ is faithful.
Under an amenability assumption this can be turned into a uniqueness result.
Corollary 5.19.
Let $\mathcal{C}$ be a left cancellative small category such that $\mathscr{G}:=\overline{\mathcal{I}_{l}}\ltimes\Omega$ is second-countable and Hausdorff. If the canonical map $C^{*}(\overline{\mathcal{I}_{l}}\ltimes\Omega)\rightarrow C^{*}_{r}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ is an isomorphism, e.g. if $\overline{\mathcal{I}_{l}}\ltimes\Omega$ is amenable, then the C*-algebra generated by a representation of $C^{*}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ that is faithful on $C^{*}(\mathcal{C}^{*}\ltimes\Omega)$ is canonically unique.
Proof.
By Theorem 5.18, a representation of $C^{*}(\overline{\mathcal{I}_{l}}\ltimes\Omega)$ that is faithful on $C^{*}(\mathcal{C}^{*}\ltimes\Omega)$ is faithful. ∎
Remark 5.20.
When the LCSC $\mathcal{C}$ is a monoid that embeds in a group $G$ the groupoid $\overline{\mathcal{I}_{l}}\ltimes\Omega$ is isomorphic to the partial transformation groupoid of a partial action of $G$ on $\Omega$, and Theorem 5.18 recovers [LS, Theorem 5.1]. Since there are left cancellative monoids that do not embed in a group and for which the groupoid $\overline{\mathcal{I}_{l}}\ltimes\Omega$ is Hausdorff, our result extends [LS, Theorem 5.1] non-trivially even for monoids. A few concrete examples are discussed in the next section.
6. Examples from integer arithmetic
The purpose of this final section is to give a concrete application of Theorem 5.18 to a monoid that does not embed in a group, so that [LS, Theorem 5.1] does not apply, and for which our ‘ideal-detecting subalgebra’ is strictly smaller than the one obtained from the interior of the isotropy as in [brown2015cartan]. The monoid arises in the study of phase transitions of C*-dynamical systems of arithmetic origin [Laca-Schulz23].
Fix $n\in\mathbb{N}^{\times}$. Then $\mathbb{N}^{\times}$ acts on $\mathbb{Z}/n\mathbb{Z}$ by right multiplication, $[k]\cdot a:=[ka]$ for $a\in\mathbb{N}^{\times}$ where $[k]=k+n\mathbb{Z}\in\mathbb{Z}/n\mathbb{Z}$ denotes the equivalence class of $k$ modulo $n$. We consider the semidirect product ×/nleft-$\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$, where the operation is defined by
$(a,[k])\cdot(b,[l])=(ab,[kb+l])\qquad a,b\in\mathbb{N}^{\times},\ k,l\in\mathbb{Z}.$
Proposition 6.1.
(cf. [Laca-Schulz23, Proposition 7.2]) The monoid ×/nleft-$\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ is left cancellative but not right cancellative, hence not embeddable in a group. Moreover,
$(a,[k])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})\cap(b,[l])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})=(\operatorname{lcm}(a,b),[0])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}),$ (6.2)
so ×/nleft-$\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ is also right LCM. The associated groupoid $\mathscr{G}$ is Hausdorff and second-countable.
Proof.
If $(ab,[kb+l])=(a,[k])(b,[l])=(a,[k])(c,[m])=(ac,[kc+m])$, then $b=c$ and $[kb+l]=[kc+m]$. This implies $[kb]=[kc]$, and hence $[l]=[m]$ in $\mathbb{Z}/n\mathbb{Z}$, hence ×/nleft-$\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ is left cancellative.
In order to see that right cancellation fails, consider for example that $(1,[2])\neq(1,[1])$, but right-multiplication by $(n,[0])$ gives $(1,[2])(n,[0])=(n,[2n+0])=(n,[0])=(n,[n+0])=(1,[1])(n,[0])$.
It is straightforward to verify (6.2) using the composition rule, see e.g. [Laca-Schulz23], proving the right LCM property. Notice that the principal right ideal $(a,[k])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$ does not depend on $[k]$. Since ×/nleft-$\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ is countable, second countability of $\mathscr{G}$ follows from Lemma 5.8 and its proof. The only thing left to show is that $\mathscr{G}$ is Hausdorff. Recall from [SpielbergSmallCats, Definition 3.2.] that a left cancellative small category $\mathcal{C}$ is finitely aligned if for every $a,b\in\mathcal{C}$ there exists a finite set $F\subseteq\mathcal{C}$ such $a\mathcal{C}\cap b\mathcal{C}=\bigcup_{c\in F}c\mathcal{C}$. Right LCM monoids are obviously finitely aligned as small categories, with $F=\{c\}$. By [LiGarsideCats, Corollary 4.2.], to prove $\mathscr{G}$ is Hausdorff it suffices to show that for all $(a,[k]),(b,[l])\in\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ the set
$X:=\big{\{}(c,[m])\in\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}\;|\;(a,[k])(c,[m])=(b,[l])(c,[m])\big{\}}$
equals $(d,[t])\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$ for some $(d,[t])\in\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$. Let $k,l,m$ be positive integers smaller than $n$, let $a,b,c\in\mathbb{N}^{\times}$ be arbitrary, and suppose that $(a,[k])(c,[m])=(b,[l])(c,[m])$. Then $(ac,[kc+m])=(bc,[lc+m])$, so we have $a=b$ because $\mathbb{N}^{\times}$ is right cancellative. Moreover $[kc+m]=[lc+m]\iff[kc]=[lc]$ so the choice of $[m]$ is arbitrary and it follows that $X=\{(c,[m])\in P\;|\;[kc]=[lc]\}$. We claim that
$X=\big{\{}(c,[m])\in\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}\;|\;[kc]=[lc]\big{\}}=\left(\frac{\operatorname{lcm}(k-l,n)}{k-l},[0]\right)\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}.$
First let $(\frac{\operatorname{lcm}(k-l,n)}{k-l}\cdot t,[p])$ be an element of the right hand side; then
$\left[k\cdot\frac{\operatorname{lcm}(k-l,n)}{k-l}\cdot t\right]-\left[l\cdot\frac{\operatorname{lcm}(k-l,n)}{k-l}\cdot t\right]=\left[\operatorname{lcm}(k-l,n)\cdot t\right]=[0].$
So $(\frac{\operatorname{lcm}(k-l,n)}{k-l}\cdot t,[p])\in X$. Conversely suppose $(c,[m])\in X$. Then $[kc]=[lc]$ implies $n$ divides $(k-l)c$ hence $\operatorname{lcm}(k-l,n)$ divides $(k-l)c$ and there is a $t$ such that $(k-l)c=\operatorname{lcm}(k-l,n)t$ hence $c=\frac{\operatorname{lcm}(k-l,n)}{k-l}t$. It follows that $(c,[m])=(\frac{\operatorname{lcm}(k-l,n)}{k-l},[0])([m],t)$ as desired. This completes the proof. ∎
Theorem 5.18 yields the following faithfulness and uniqueness result.
Corollary 6.3.
A representation of $C^{*}(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$ is faithful if and only if its restriction to the crossed product $C(\Omega)\rtimes(\{1\}\ltimes\mathbb{Z}/n\mathbb{Z})$ is faithful.
Proof.
It is easy to verify that (×$(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}=\{1\}\rtimes\mathbb{Z}/n\mathbb{Z}$. The result now follows from Theorem 5.18, whose assumptions are satisfied by Proposition 6.1 and Proposition 5.17. ∎
Next we aim to show that the conditions for detecting ideals from [brown2015cartan] and our generalization to that from [LS] are effectively different. In order to do this we first need to compute the interior of the isotropy.
Proposition 6.4.
Let $S\curvearrowright\Omega$ be an action of an inverse semigroup on a topological space. The interior of the isotropy of the action groupoid $\mathscr{G}:=S\ltimes\Omega$ consists of all pairs $[s,x]\in\mathscr{G}$ such that there exists an open neighborhood $U$ of $x$ for which $s\cdot y=y$ for all $y\in U$.
Proof.
A basis for the topology of $S\ltimes\Omega$ is given by the subsets $[s,U]=\{[s,\chi]\;|\;\chi\in U\}$ for $s\in S$ and $U$ an open subset of $D_{s^{-1}s}$. It follows that a point $[s,x]\in S\ltimes\Omega$ is in the interior of the isotropy if and only if we can choose $U\subseteq D_{s^{-1}s}$ such that $s$ fixes every element of $U$. ∎
Proposition 6.5.
Let $\mathcal{C}$ be a left cancellative small category. The interior of the isotropy of $\mathscr{G}:=\overline{I_{l}}\ltimes\Omega$ is the collection of all pairs $[s,\chi]$ such that there exists an idempotent $e\in E(\overline{\mathcal{I}_{l}})$ satisfying
  • $e\leq s^{-1}s$
  • $\chi(e)=1$
  • for each $c\in\operatorname{dom}(e)\subseteq\mathcal{C}$ there exists $u\in\mathcal{C}^{*}$ such that $[s,\chi_{c}]=[cuc^{-1},\chi_{c}]$.
Proof.
Let $[s,\chi]\in\operatorname{Iso}(\mathscr{G})^{\circ}$; then there is a basic open neighborhood $\Omega(e)=\{\psi\in\Omega\;|\;\psi(e)=1\}$ of $\chi$ such that $s\cdot\psi=\psi$ for all $\psi\in\Omega(e)$. In particular, if $c\in\operatorname{dom}(e)$, then $\chi_{c}\in\Omega(e)$ and $s$ fixes $\chi_{c}$ so $[s,\chi_{c}]=[cuc^{-1},\chi_{c}]$ for some $u\in\mathcal{C}^{*}$. Conversely, if $[s,\chi]\in\mathscr{G}$ and there is an idempotent $e\leq s^{-1}s$ such that $\chi(e)=1$ and $[s,\chi_{c}]\in[c\mathcal{C}^{*}c^{-1},\chi_{c}]$, then $s$ fixes a dense subset of $\Omega(e)$ so $s$ fixes every $\psi\in\Omega(e)$ by continuity. ∎
Next we show that the action groupoid of invertible elements acting on the unit space is a more efficient way to detect ideals of $\mathcal{T}_{\lambda}(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$ than the interior of the isotropy.
Proposition 6.6.
(×$(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}\ltimes\Omega\subsetneq\operatorname{Iso}(\mathscr{G})^{\circ}$.
Proof.
First we show that (×$(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}\ltimes\Omega\subset\operatorname{Iso}(\mathscr{G})^{\circ}$. Suppose $(1,[k])\in(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}$ and let $(a,[m])\in\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$. Then $(1,[k])\cdot\chi_{(a,[m])}=\chi_{(a,[ka+m])}=\chi_{(a,[m])}$ because $(a,[ka+m])=(a,[m])$ in ×/nleft-$\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}$. Since the characters $\chi_{(a,[m])}$ are dense in $\Omega$ by Lemma 5.8, we see that (×$(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}\ltimes\Omega$ is contained in the isotropy at every point. But (×$(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}\ltimes\Omega$ is open, hence it is actually contained in the interior of the isotropy.
Next define an element $s$ of the left inverse hull by
$s:=(n,[0])(1,[1])(n,[0])^{-1}:(n,[0])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$xlongrightarrow$\xlongrightarrow{\cong}(n,[0])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$
$(n,[0])(a,[m])$ $\longmapsto(n,[0])(a,[a+m]).$
We will show that
$[s,\chi_{(n,[0])}]\in\operatorname{Iso}({(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})\ltimes\Omega_{2}})^{\circ}\setminus(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}\ltimes\Omega.$
We show first that the element $[s,\chi_{(n,[0])}]$ is in the interior of the isotropy of $\mathscr{G}$. First note that points of the form $\chi_{(n,[0])(b,[m])}=\chi_{(nb,[m])}$, with $b\in\mathbb{N}^{\times}$ and $10\leq m\leq n-$ 1, are dense in $V:=\Omega((n,[0])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z}))$. For every such point we have $s\cdot\chi_{(nb,[m])}=\chi_{(nb,[m+b])}=\chi_{(nb,[m])}$ where the last equality is due to the fact that for each $e\in E(\overline{\mathcal{I}_{l}})$ we have $(nb,[m])\in\operatorname{dom}(e)\iff(nb,[k])\in\operatorname{dom}(e)$ for all $10\leq k\leq m-$ 1. Thus $s$ fixes a dense set of points in $V$ and must fix every point in $V$ by continuity. Then $[s,\chi_{(n,[0])}]\in\operatorname{Iso}(\mathscr{G})^{\circ}$ by Proposition 6.4.
To conclude that $[s,\chi_{(n,[0])}]\notin(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}\ltimes\Omega$ we need to show that for all $e\leq s^{-1}s$ such that $(n,[0])\in\operatorname{dom}(e)$ and $u=(1,[k])\in(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})^{*}$ we have $se\neq ue$. Suppose for contradiction that $e\in E$ and $u=(1,[k])$ are as above and that $se=ue$ and let $(na,[m])\in\operatorname{dom}(e)\subseteq(n,[0])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$ be arbitrary. Then $se=ue$ implies that $(na,[m+a])=s(na,[m])=(1,[k])(na,[m])=(na,[m])$ so $[m+a]=[m]$ and $n$ must divide $a$. But then $\operatorname{dom}(e)\subseteq(n^{2},[0])(\mathbb{N}^{\times}\ltimes\mathbb{Z}/n\mathbb{Z})$ which contradicts the assumption that $(n,[0])\in\operatorname{dom}(e)$. ∎
Remark 6.7.
A similar result to Proposition 6.6 can be proved about the monoid $\mathbb{N}^{\times}\ltimes\mathbb{Z}$, which is group embeddable, and $\mathbb{N}^{\times}\ltimes\mathbb{Q}/\mathbb{Z}$, which is not.
In contrast to the above applications we now consider the submonoid $\mathbb{Z}\rtimes\mathbb{N}^{\times}$ of the group $\mathbb{Q}\rtimes\mathbb{Q}_{+}^{*}$, where the operation is given by $(k,a)(l,b)=(k+al,ab)$ and the set of invertible elements is (×)*=$(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}=\mathbb{Z}\rtimes\{1\}$. This monoid is the opposite of $\mathbb{N}^{\times}\ltimes\mathbb{Z}$ and corresponds to the additive boundary quotient in [LacaRaeburn]. It has a finer principal ideal structure, see [LacaRaeburn, Section 5] and the isotropy at principal characters is consequently smaller. As a result, we see that ideal detection is more efficient using the interior of the isotropy, which in this case is simply the unit space.
Proposition 6.8.
Let $\Omega:=\mathscr{G}^{(0)}$ be the unit space of the groupoid $\mathscr{G}$ associated to the monoid $\mathbb{Z}\rtimes\mathbb{N}^{\times}$. Then $\operatorname{Iso}(\mathscr{G})^{\circ}\subsetneq(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}\ltimes\Omega$.
Proof.
The action of $(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}$ on $\Omega$ is topologically free by [ToeLaNe, Lemma 1.6], and by [LS, Theorem 5.9] so is the partial action of $\mathbb{Q}_{+}^{*}\rtimes\mathbb{Q}$. Thus the interior of the isotropy is the unit space, which is properly contained in the action subgroupoid (×)*$(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}\ltimes\Omega$. For example, the element $((1,1),\chi_{(0,2)})\in(\mathbb{Z}\rtimes\mathbb{N}^{\times})^{*}\ltimes\Omega$ is not in $\operatorname{Iso}(\mathscr{G})$ because $(1,1)\cdot\nolinebreak\chi_{(0,2)}=\chi_{(1,2)}\neq\chi_{(0,2)}$. ∎